OpenStudy (anonymous):

Find the product. -2x a(-4x b - 2x 3 + 5x) -8xa+b - 4xa+3 - 10xa+1 8xa+b + 4x3a - 10xa 8xa+b + 4xa+3 - 10xa+1

OpenStudy (anonymous):

(−2x)(a)((−4x)(b)−2x(3)+5x) =((−2x)(a))((−4x)(b)+−6x+5x) =((−2x)(a))((−4x)(b))+((−2x)(a))(−6x)+((−2x)(a))(5x) =8abx2+12ax2−10ax2 =8abx2+2ax2